Make New Entry,
Make Followup Entry
User name septum_auto
Log entry time 20:33:22 on May 2, 2012
Entry number 381534
keyword=Wrong septum current, please check!
Left septum set current 50A is not correct, should be 361.847A! please check!
Left septum read back current 49.2741A is not correct, should be 50A! please check!
Right septum set current 50A is not correct, should be 474.699A! please check!
septum alarm using halog_post [halog_post 1.1]
A copy of this log entry has been emailed to: email@example.com